xcostheta+ysintheta=c and xcostheta-ysintheta=d elininate theta

Grade:11

1 Answers

Shalini
6 years ago
let theta be A thenXcosA+YsinA=CXcosA-YsinA=D1.Adding equation (1 and 2)2XcosA=C+DCosA=C+D/2XA=Cos^-1(C+D/2X)2.subtracting equation (1 and 2)2YsinA=C-DSinA=C-D/2YA=sin^-1(C-D/2Y)